Shift report 20030218 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Feb 18 2003 - 23:54:06 EST

  • Next message: Apache: "Shift report 20030219 00:00-08:00"
    ------------------------------------------------------------
    Date: Tue Feb 18 23:54:06 2003, Shift: 16:00-24:00
    Supervisor: SZ, with: 
    Log book#14 pages 24 - 25
    
    Shift Summary:
    ------------------------------------------------------------
    1. Machine Status
       ---------------  
       The beam stayed cogged  from the begginning of shift.
       It has been dumped at ~ 22:00. 
       Once in a while the background  was  very high
    --------------------------------------------------
    2. Achievements
       ------------
       I proceed to take the last run from the previous shift.
       Data Run : MRS 40deg, 1050A, FS 12deg, 1/4B
                                               
       Run 8011 :  85K trigger3, 17K trigger6 
       Run 8012 : 125K trigger3, 28K trigger6
       Run 8013 :  26K trigger3,  3K trigger6
       Run 8014 :  64K trigger3, 18K trigger6; 
    
       Trigger4 scale down set to 1000 (from 500), and
       Trigger3 scale down set to 2(from 1) for Runs 8011,    8012,8014
    
    --------------------------------------------------------
    3. Detector Problems
       ------------------
     16:46  DVM and Si detectors tripped! L1458A was turned off itself!
     17:18  TPM1 and TPM2 tripped. 
     17:24-  T1 and T2 tripped, also. 
    
    --------------------------------------------------------
    Plan :
      
      MRS : need  50K trigger3 at 40-degree, 1050B 
       FS : need 100K trigger6 at 12-degree  1/4B
    
    
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Tue Feb 18 2003 - 23:58:27 EST